OpenStudy (anonymous):

whats the equation for: a. (b+c) if a = 5, b= 8 and c= 4?

OpenStudy (anonymous):

this is a dotvector so its a.(b+c) (aDOT(b+C)

OpenStudy (anonymous):

here is a visual of the question: its iii

OpenStudy (anonymous):

a=5,b=8,c=4 b+c=12,suppose,b+c=B and a=5=A a.(b+c)=A.B=ABcostheta

OpenStudy (anonymous):

is b+ c still 12 if you look at the attachment above?

OpenStudy (anonymous):

well,it should be ,b+c=(Bx+Cx)i+(By+Cy)j+(Bz+Cz)k i,j,k=unit vector

OpenStudy (anonymous):

ok now im confused whare do i find x es and ys? and the i j k? if you dont mind can you plsease sub in the values?

OpenStudy (anonymous):

but here,b+c=8

OpenStudy (anonymous):

you don't have to,since they have given the exact value of b and c

OpenStudy (anonymous):

i see so i shuld to the onther questions in the attachment the same way?

OpenStudy (anonymous):

like the first questions is 8?

OpenStudy (anonymous):

and the second is 80 degrees?

OpenStudy (anonymous):

suppose,a=A,b+c=B so,a.(b+c)=A.B=ABcostheta I suppose